# Motivic cohomology of cyclic coverings

### v2

@@ 4.4 Bicyclic coverings of Koras-Russell threefolds of the first kind
with $f=-y\in\mathcal{O}_{X}(X)$ of orders $\alpha,\beta$. The discussion of Section 4.1 and Theorem 3.32 imply that $CH^{i}(Y_{\alpha,\beta})=0$ for $i\geq 1$. It follows that all vector bundles over $Y$ are stably trivial; in fact, it follows that at least all vector bundles of rank $\neq 2$ are trivial by well-known cancellation theorems (cf. [FRS], [S2]). All vector bundles of rank $2$ are trivial if and only if the Hermitian $K$-theory group $W_{SL}(Y_{\alpha,\beta})$ is trivial; in fact, the vanishing of $H_{Nis}^{2}(Y_{\alpha,\beta},\textbf{I}^{3})$ would already guarantee the triviality of all vector bundles over $Y_{\alpha,\beta}$ (cf. [Sy]). Theorem 3.34 implies that all vector bundles over $Y_{\alpha,\beta}$ are indeed trivial as soon as $\alpha$ and $\beta$ are both odd integers. Finally, if $k=\mathbb{C}$, then $Y_{s,t}$ is topologically contractible by Remark 3.33.

### v3

@@ 4.4 Bicyclic coverings of Koras-Russell threefolds of the first kind
with $f=-y\in\mathcal{O}_{X}(X)$ of orders $\alpha,\beta$. The discussion of Section 4.1 and Theorem 3.30 imply that $CH^{i}(Y_{\alpha,\beta})=0$ for $i\geq 1$. It follows that all vector bundles over $Y$ are stably trivial; in fact, it follows that at least all vector bundles of rank $\neq 2$ are trivial by well-known cancellation theorems (cf. [FRS], [S2]). All vector bundles of rank $2$ are trivial if and only if the Hermitian $K$-theory group $W_{SL}(Y_{\alpha,\beta})$ is trivial; in fact, the vanishing of $H_{Nis}^{2}(Y_{\alpha,\beta},\textbf{I}^{3})$ already guarantees the triviality of all vector bundles over $Y_{\alpha,\beta}$ (cf. [Sy]). Finally, Theorem 3.33 implies that all vector bundles over $Y_{\alpha,\beta}$ are indeed trivial. If $k=\mathbb{C}$, then $Y_{s,t}$ is topologically contractible by Remark 3.31.